OpenStudy (anonymous):

A model scale is 1 cm = 8 m. If the actual object is 40 meters, the model is_______ cm long.

OpenStudy (anonymous):

this is a propertions question: what is is asking is can you find X: 1cm/8m = Xcm/40m

OpenStudy (anonymous):

so if you cross-multiply, you get 40 (1) = X (8) X= 40 (1)/8 = 5 the model will be 5 centimetres long

OpenStudy (anonymous):

yes use ratio and proportion

OpenStudy (anonymous):

x cm 1 cm ---- = ------ solve for x here to get 40 m 8 m x= 40(1)/8 = 5cm ans...